How do I solve by elimination: X-4Y=-23
2X+3Y=9

Answer by Annabelle1(69)

we need to get the coefficient of x to be the same.
multiply the first equation through by 2
2x-8y=-46
2x+3y=9
subtract the two equations to eliminate the x
2x-2x-8y-3y=-46-9
-11y=-55
y=-55/-11
y=5
sub this number back into one of the original equations to get your x value
2x-8*5=-46
2x-40=-46
2x=-46+40
2x=-6
x=-3
your solution is x=-3, y=5